OpenStudy (anonymous):

solve the following system of equations.-3x + 5y=68, x + 4y=51

OpenStudy (anonymous):

x+4y=51 x=51-4y then, substitute it to -3x

OpenStudy (anonymous):

-3x+5y=68 -3(51-4y)+5y=68 -153+12y+5y=68 17y=221 y=13 now using y u have got, substitute it to x+4y=51 and u would have x

OpenStudy (anonymous):

so the answer would finally be?